Week of 170619

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday - virtual meeting

  • Note: Mon-Thu this week the WLCG Workshop is being held in Manchester.

  • You may provide relevant incidents, announcements etc. for the operations record.

Attendance:

  • local:
  • remote:

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Mostly MC production and group production for analysis, running at 250-300k job slots level, data rebalancing between sites and normal transfer activities
      • Atlas Software and Computing week
    • Issues:
      • Quiet week apart from the
      • DNS network problem at CERN on Sunday evening
        • induced T0 jobs failing because no access to the conditions DB
        • Systems recovered quickly when network came back, but central service Kibana monitoring (no easy way to check that ATLAS central services are ok).
      • CVMFS monitoring issue: ASGC stratum 1 was removed from atlas-condb.cern.ch and atlas.cern.ch monitoring pages (due to connectivity problems, the TW cvmfs stratum 1 was removed from the official list of WLCG stratum 1 after discussion in the February GDB) => this solved the CVMFS monitoring instabilities; however still some replica loading for ever on the monitoring page.

  • CMS reports ( raw view) -
    • NTR request for resources from the physics community still on the low side (Stefano)

  • ALICE -
    • continued very high activity
    • some fallout from the DNS incident at CERN on Sunday (OTG:0038195)

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF:
  • EGI:
  • FNAL:
  • IN2P3:
  • JINR: Configuration of IPv6 on SEs completed. Some glitches on the main router related to IPv6 on 05-05-2017 & 08-05-2017, solved.
  • KISTI:
  • KIT:
  • NDGF:
  • NL-T1:
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL: Increase OPN link from 2 * 10Gbit links to 3 * 10Gbit links delayed until Wednesday 28th June.
  • TRIUMF:NTR

  • CERN computing services: Several CERN services, notably CVMFS, were affected by a hypervisor outage on the 13th of June. Please refer to the SIR: ServiceIncidentCloud20170613 for further details. On Sunday 18th of June, a network problem affecting the DNS servers also lead to degraded services at CERN. (Ref. OTG:0038195)
  • CERN storage services:
    • EOS instances degraded during the DNS servers overload.
    • EOSALICE didn't recover by itself, one of the daemon had to be restarted.
  • CERN databases:
  • GGUS:
  • Monitoring:
  • MW Officer: NTR
  • Networks:
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2017-06-24 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback